OpenStudy (anonymous):

taylor series of f(x)=x-x^3, a=-2 with radius of convergence and interval of convergence

OpenStudy (amistre64):

define f' f'' f''' and such tell it goes to zero

OpenStudy (amistre64):

f = x - x^3 f' = -3x^2 f'' = -6x f''' = -6 f'''' and beyond = 0

OpenStudy (anonymous):

yes i got that far

OpenStudy (amistre64):

define the f's at=a=-2

OpenStudy (anonymous):

6, -11, 12, and ,6

OpenStudy (amistre64):

f = -2 - (-2)^3 = 6 f' = -3(-2)^2 = -12 f'' = -6(-2) = 12 f''' = -6 = -6 f'''' and beyond = 0 = 0

OpenStudy (amistre64):

i did forget a 1 didnt i

OpenStudy (amistre64):

6,-11,12,-6, 000000

OpenStudy (anonymous):

i can't find a pattern for the series

OpenStudy (amistre64):

set these up into the standard taylor getup as the coeefs \[x-x^3=6\frac{x^0}{0!}-11\frac{x^1}{1!}+12\frac{x^2}{2!}-6\frac{x^3}{3!}+0000\]

OpenStudy (amistre64):

\[x-x^3=6-11(x+2)+6(x+2)^2-(x+2)^3\]

OpenStudy (anonymous):

I'm stuck right there!

OpenStudy (amistre64):

lol, umm, just expand it out and you should see that it siplifies to x-x^3

OpenStudy (amistre64):

the polynomial expansion of a polynomial expansion ... is just the polynomial you started with :)

OpenStudy (anonymous):

so I don't need to put it in the form of \[\sum_{n=0}^{\infty} (x-a)^n\] etc etc?

OpenStudy (amistre64):

no. polynomials are unique; therefore you cant really have 2 different polynomials of the same curve. the exercise is meant to get you familiar with the concept of createing a power series using something simple

OpenStudy (amistre64):

in other words, it means to solidify techniques

OpenStudy (anonymous):

so how would I find out the radius and interval of convergence?

OpenStudy (amistre64):

the radius of convergence is the distance from "a" wherein the taylor matches the function since the taylor matches the function at all point; the radius in infinity

OpenStudy (amistre64):

the interval is the range that the radius spans; -R < x < R ; and since R = inf on this .. interval is just (-inf,inf)

OpenStudy (amistre64):

dont live life in a formula :) understand the concepts

OpenStudy (anonymous):

okay thanks again!!!

OpenStudy (amistre64):

yep